OpenStudy (anonymous):

ok on the calculator when i punch in -2 to the power of 4 i get -16 and when i punch in (-2) to the power of 4 i get 16 why ? can someone please explain

OpenStudy (saifoo.khan):

There is a difference, -2^4 -1(2*2*2*2) -1(16) -16 in (-2^4) -2*-2*-2*-2 16

OpenStudy (anonymous):

the first one doesnt include the negative sign, its just adding a negative sign to 2 to the power of 4 the other way, its the whole number, -2, to the power of 4

OpenStudy (aroub):

lol i love how u punch it in.. (-2)^4 when ever u have a even exponent the number becomes positive (2)^4 this is like saying 2*2*2*2 = 16 or 4*4=16 -2^4 = -2*-2*-2*-2= -16

OpenStudy (anonymous):

thanks